OpenStudy (bookworm14):

Can ya'll please check this for me ?!?!?! I will have to post the equation (s) below

OpenStudy (bookworm14):

\[\left| 3-9x \right| \le 39\] I got\[x \le -4 ~~or ~~x \ge -4 2/3\]

OpenStudy (bookworm14):

that is supposed to be -4 and 2/3

OpenStudy (anonymous):

|3-9x|<=39 this implies -39<=3-9x<=39 divide by 3 then we have -13<=1-3x<=13 This implies -14<=-3x<=12 Divie by -3 and the inequality changes direction this implies 14/3>=x>=-4.@Bookworm14

OpenStudy (anonymous):

@Bookworm14